Okay here i go again with another written problem. Please help.
The measure of one angle of a triangle is twice the measure of a second angle. The measure of the third angle is 12 less than the sum of the other two. Find the measures of the angles of the triangle.

Answer by ankor@dixie-net.com(22740)

Let the second angle = x
:
The measure of one angle of a triangle is twice the measure of a second angle.
This says: 1st angle = 2x
:
The measure of the third angle is 12 less than the sum of the other two.
This says: 3rd angle = (x + 2x) - 12 = (3x-12)
:
Find the measures of the angles of the triangle.
:
2nd angle + 1st angle + 3rd angle = 180 (like all triangles)
x + 2x + (3x - 12) = 180
6x = 180 + 12
x = 192/6
x = 32 degree is the 2nd angle
:
1st angle = 2(32) = 64
:
3rd angle = 32 + 64 - 12 = 84 degrees
:
Check: 32 + 64 + 84 = 180